# LargeAD: Large-Scale Cross-Sensor Data Pretraining for Autonomous Driving

### v2

@@ Abstract
Index Terms:
Autonomous Driving; 3D Scene Understanding; LiDAR Segmentation; Representation Learning; Data Pretraining
@@ 1 Introduction
The emergence of large language models (LLMs) [1, 2, 3, 4] has transformed natural language processing, paving the way for similar breakthroughs in computer vision through vision foundation models (VFMs) such as SAM [5], X-Decoder [6], and SEEM [7]. These models have demonstrated remarkable capabilities in extracting rich pixel-level semantics from 2D images. However, extending these advancements to the 3D domain remains a largely unexplored frontier. As autonomous driving applications increasingly rely on 3D data from LiDAR sensors, there is a growing need to transfer the success of VFMs in 2D vision to 3D scene understanding.
Accurate perception of LiDAR point clouds is crucial for safe autonomous driving and advanced driver-assistance systems [8, 9, 10, 11, 12, 13]. Traditional LiDAR point cloud models often depend on large annotated datasets, which are costly and time-consuming to create [14, 15]. To alleviate this challenge, research has explored semi-supervised [16, 17] and weakly-supervised [15, 18] methods. However, these approaches suffer from limited generalizability, especially when faced with diverse sensor configurations, e.g., different LiDAR beam numbers, camera placements, sampling rates, and potential sensor corruptions [19, 20, 9, 21, 22, 23, 24, 25, 26, 27]. This limitation poses a significant challenge for real-world scalability.
@@
A core innovation in our framework is the use of VFMs to generate semantically enriched superpixels from camera images, which are then aligned with LiDAR data to construct high-quality contrastive samples, as shown in Fig. 1. These semantic superpixels provide enhanced 2D-3D correspondences that capture object-level coherence, reducing the errors commonly associated with over-segmentation and “self-conflict” in contrastive learning [29]. This alignment significantly improves performance in downstream tasks, including 3D object detection and segmentation.
@@
As shown in Fig. 2, compared to state-of-the-art methods such as SLidR [30] and ST-SLidR [31], our framework introduces significant advancements: i) the use of semantically rich superpixels to resolve the “self-conflict” issue in contrastive learning, ii) the creation of high-quality contrastive samples, which lead to faster and more stable convergence, and iii) reduced computational overhead due to the more efficient superpixel generation process.
In summary, this paper offers the following contributions:
  • We introduce LargeAD, a scalable, consistent, and generalizable framework designed for large-scale pretraining on data captured from onboard vehicle sensors, addressing the challenges of diverse LiDAR configurations and improving representation learning capabilities.
  • To the best of our knowledge, this is the first comprehensive study to explore pretraining across multiple large-scale driving datasets, leveraging cross-dataset knowledge to enhance the model’s generalizability to different sensor setups and driving environments.
  • Our framework incorporates several key innovations: i) VFM-based superpixel generation that enriches semantic representation, ii) VFM-assisted contrastive learning to align 2D-3D features, iii) superpoint temporal consistency to stabilize point cloud representations across time, and iv) multi-source data pretraining that ensures robustness across domains.
  • Our approach demonstrates significant performance advantages over state-of-the-art methods, achieving superior results in both linear probing and fine-tuning tasks across $11$ diverse point cloud datasets, showcasing its adaptability and efficiency in real-world applications.
The remainder of this paper is organized as follows. Sec. 2 reviews the relevant literature on autonomous driving data perception and pretraining, as well as multi-dataset utilization. Sec. 3 elaborates on preliminaries of image-to-LiDAR contrastive learning. Sec. 4 presents the technical methodology behind the proposed large-scale cross-sensor pretraining framework. Finally, Sec. 5 presents the experimental validation of our approach, and Sec. 6 concludes the paper with discussions on future directions.
@@ 2 Related Work
In this section, we provide a thorough literature review of relevant works in the areas of LiDAR-based autonomous driving scene understanding (Sec. 2.1), popular vision foundation models (Sec. 2.2), different 3D representation learning techniques (Sec. 2.3), and multi-dataset learning and utilization in the domain of 3D object detection and LiDAR semantic segmentation (Sec. 2.4).
@@ 2.1 LiDAR-Based Scene Understanding
For autonomous vehicles, accurate and dense 3D perception is crucial for safe navigation [10, 33, 34]. Researchers have developed various methods for point cloud segmentation, including those based on raw points [35, 36, 37, 38, 39], range views [40, 41, 42, 43, 44, 45, 46], bird’s eye views [47, 48], voxels [49, 50, 51, 52], and multi-view fusion [53, 54, 55, 56, 57, 58]. Despite achieving significant advancements, these models typically rely on extensively annotated datasets, posing scalability issues [19]. To alleviate the annotation burden, recent studies have explored semi-supervised [16, 17, 59], weakly-supervised [15, 60, 61, 62, 18], and active learning [63, 64, 65] methods, as well as domain adaptation techniques [66, 67, 68, 69, 70, 71]. This work utilizes a self-supervised learning strategy, distilling knowledge from VFMs through camera-to-LiDAR associations, thereby eliminating the need for manual annotations during pretraining.
@@ 2.4 Multi-Dataset Utilization
Leveraging multiple datasets has emerged as a promising approach for improving the generalization of LiDAR-based models in autonomous driving [109]. Recent works like MDT3D [110] and MS3D++ [111] explored multi-source training for 3D object detection, while addressing challenges such as label space conflicts. Similarly, methods like M3Net [112] utilized unified label spaces for semantic segmentation, demonstrating the advantages of multi-dataset learning. In this work, we extend these efforts by proposing a multi-source pretraining strategy that incorporates data from diverse LiDAR datasets, each with unique sensor setups and conditions [113, 114, 115, 12, 116]. By aligning 2D and 3D features across different sources, our approach improves cross-modal consistency and robustness. To the best of our knowledge, this is the first comprehensive study on multi-source pretraining for image-to-LiDAR representation learning, allowing our framework to generalize more effectively across varied driving environments and sensor configurations.
@@ 3 Image-to-LiDAR Data Pretraining
In this section, we elaborate on the technical details of the image-to-LiDAR pretraining. We first present the preliminaries (Sec. 3.1), followed by the introduction of superpixel-driven contrastive learning (Sec. 3.2), and conclude with the description of the contrastive learning objective (Sec. 3.3).
@@ 3.1 Problem Formulation
$[u_{i},v_{i},1]^{\text{T}}=\frac{1}{z_{i}}\times\Gamma_{K}\times\Gamma_{c\leftarrow l}\times[x_{i},y_{i},z_{i}]^{\text{T}}~,$ (1)
@@ 4 LargeAD: A Scalable, Versatile, and Generalizable Framework
In this section, we present the technical details of the proposed framework. We begin by introducing our VFM-based superpixel generation approach (Sec. 4.1), followed by the explanation of semantic spatial consistency learning (Sec. 4.2). Next, we discuss the instance superpoint temporal consistency (Sec. 4.3), and conclude with the description of multi-source data pretraining (Sec. 4.4), culminating with a summary of the overall framework (Sec. 4.5).
@@ 4.1 Superpixel Generation from Foundation Models
Previous works have utilized SLIC [32] to group visually similar image regions into superpixels. However, SLIC often over-segments semantically coherent areas (see Fig.2), which poses challenges for contrastive learning, particularly due to the “self-conflict” phenomenon. This occurs when semantically similar superpixels are incorrectly treated as negative samples[120]. While [31] introduced a semantically tolerant loss to address this issue, the lack of high-level semantic understanding in SLIC exacerbates the difficulties in contrastive learning. To overcome these challenges, we generate semantic superpixels using vision foundation models (VFMs), which provide semantically rich superpixels and significantly improve the representation learning for both near and far points in the LiDAR point cloud (see Fig. 5).
Instead of relying on low-level RGB features, our approach enhances superpixel generation by leveraging VFMs derived from large-scale pretrained image encoders [5, 7, 28, 6]. Unlike SLIC, VFMs capture high-level semantic information (as shown in Fig. 2), allowing us to create more semantically meaningful superpixel sets, denoted as $\widehat{\Phi}_{\mathcal{S}}=\left\{\left\{\mathbf{s}_{m_{v}}\mid m_{v}=1,\ldots,M_{v}\right\}\mid M_{v}\ll M\right\}$. The generation process begins with the creation of semantic masks via prompts. By incorporating more abstract features, VFMs effectively address the “self-conflict” issue by grouping semantically similar regions more coherently, reducing the risk of misclassification during contrastive learning. As a result, the generated superpixels more accurately represent object semantics rather than just visual similarities. The corresponding superpoint set, $\widehat{\Phi}_{\mathcal{O}}=\left\{\left\{\mathbf{o}_{m_{v}}\mid m_{v}=1,\ldots,M_{v}\right\}\mid M_{v}\ll M\right\}$, is established using Eq. 1, ensuring proper alignment between the 2D image features and the 3D LiDAR point features.
@@ 4.2 Semantic Spatial Consistency Learning
To implement this framework, we use the same trainable LiDAR point cloud encoder $F_{\theta_{p}}$ and frozen image encoder $G_{\theta_{i}}$ as described earlier, extracting features from the LiDAR point cloud and 2D images, respectively. For the contrastive loss, we employ the projection heads $H_{\omega_{p}}$ and $H_{\omega_{i}}$ from Sec. 3.2, projecting both point and image features into a shared DD–dimensional embedding space. Unlike the low-level cues generated by SLIC, the superpixels produced by VFMs are enriched with semantic information, leading to more coherent and meaningful representations.
@@ 4.3 Instance Superpoint Temporal Consistency
Since the target embedding for all points within a segment in the current frame serves as the mean segment representation from the next frame, this loss will force points from a segment to converge to a mean representation while separating from other segments, implicitly clustering together points from the same instance. Fig. 4 provides the positive feature correspondence in our contrastive learning framework. Furthermore, we swap $\widehat{\mathcal{F}}^{t}$ when generating the target mean embedding features to form a symmetric representation. In this way, the correspondence is encouraged from both $t\to{t+1}$ and ${t+1}\to t$, which leads to the following optimization objective: ${\mathcal{L}}^{\mathrm{tmp}}={\mathcal{L}}^{t\to{t+1}}+{\mathcal{L}}^{{t+1}\to t}$.
@@ 4.4 Multi-Source Data Pretraining
$\mathcal{L}^{\mathrm{cdp}}=-\frac{1}{M}{\sum_{i=1}^{M}\log\left[\frac{e^{\left(\left<\mathbf{k}^{(m)}_{i},\mathbf{k}^{(n)}_{i}\right>/\tau\right)}}{\sum_{j=1}^{M}e^{\left(\left<\mathbf{k}^{(m)}_{i},\mathbf{k}^{(n)}_{j}\right>/\tau\right)}}\right]}~.$ (7)
@@
The multi-source data pretraining leverages a variety of data sources to create a more resilient and flexible model. By addressing the substantial distributional differences between domains, feature normalization ensures consistency across diverse datasets, facilitating a more unified representation space. The cross-source contrastive loss further enhances adaptability, enabling it to transfer knowledge across multiple domains while maintaining domain-specific distinctions. As we will show in Sec. 5, this strategy results in improved generalization across different LiDAR sensors, leading to stronger performance on out-of-source downstream tasks.
TABLE I: Benchmark study of data pretraining methods that are pretrained on nuScenes [123] and fine-tuned on nuScenes [113], SemanticKITTI [114], and Waymo Open [115]. The evaluations are conducted on the official validation split of each dataset. Group (a): random initialization. Group (b): single-modality data pretraining. Group (c): image-to-LiDAR data pretraining by distilling ResNet [124]. Groups (d), (e), and (f): image-to-LiDAR data pretraining by distilling ViT [125] (Small, Medium, Large) from DINOv2 [74]. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores under each group are highlighted in bold and underline.
@@ 4.5 Overall Framework
$\mathcal{L}=\mathcal{L}^{\mathrm{vfm}}+\mathcal{L}^{\mathrm{tmp}}+\mathcal{L}^{\mathrm{p2s}}+\mathcal{L}^{\mathrm{cdp}}~.$ (8)
@@ 5 Experiments
In this section, we conduct extensive experiments to validate the effectiveness of the proposed LargeAD framework. We first introduce the datasets (Sec. 5.1) and describe the implementation details (Sec. 5.2). Next, we present a comparative study, demonstrating the superior performance of our framework compared to state-of-the-art methods across multiple tasks and datasets (Sec. 5.3). Finally, we provide an in-depth ablation study to analyze the impact of different components and data sources on the performance (Sec. 5.4).
@@ 5.1 Datasets
TABLE II: Benchmark study of data pretraining methods that are: (a) random initialized, (b) pretrained on SemanticKITTI [114], (c) pretrained on Waymo Open [115], and (d) pretrained based on the proposed large-scale cross-sensor data pretraining pipeline. After pretraining, all methods are fine-tuned on SemanticKITTI [114], nuScenes [113], and Waymo Open [115], respectively, and evaluated on the official validation split of each dataset. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
TABLE III: Comparisons of different data pretraining methods that are: (a) random initialized, (b) pretrained on nuScenes [123], (c) pretrained on SemanticKITTI [114], (d) pretrained on Waymo Open [115], and (e) pretrained based on the proposed large-scale cross-sensor data pretraining pipeline. After pretraining, all methods are fine-tuned on different downstream datasets [15, 130, 131, 132, 133, 134, 70], respectively, and evaluated on the official validation split of each dataset. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
Figure 7: The qualitative results of object detection trained with $5\%$ labeled data. The first row shows the model trained with random initialization, while the second row displays results from our proposed framework. The groundtruth / predicted results are highlighted with blue / red boxes, respectively. Best viewed in colors and zoomed-in for additional details.
TABLE IV: Robustness probing of data pretraining methods under eight out-of-distribution corruptions in the nuScenes-C dataset from the Robo3D benchmark [135]. The mCE score is the lower the better while mRR and mIoU scores are the higher the better. All mCE, mRR, and mIoU scores are given in percentage (%). Avg denotes the average mIoU scores of methods across all eight corruptions. The best and second best scores are highlighted in bold and underline.
@@
TABLE V: Comparisons of state-of-the-art pretraining methods pretrained and fine-tuned on nuScenes [113] using the specified data proportions. All methods employ CenterPoint [136] or SECOND [137] as the 3D object detection backbones. All sores are given in percentage (%). The best and second best scores are highlighted in bold and underline.

### v3

# ] [Uncaptioned image] WorldBench Team

@@ Abstract
[…]
[ [Uncaptioned image] Project Page]https://ldkong.com/LargeAD
@@ 1 Introduction
The emergence of large language models (LLMs) [1, 2, 3, 4, 5] has transformed natural language processing, paving the way for similar breakthroughs in computer vision through vision foundation models (VFMs) such as SAM [6], X-Decoder [7], and SEEM [8]. These models have demonstrated remarkable capabilities in extracting rich pixel-level semantics from 2D images. However, extending these advancements to the 3D domain remains a largely unexplored frontier. As autonomous driving applications increasingly rely on 3D data from LiDAR sensors, there is a growing need to transfer the success of VFMs in 2D vision to 3D scene understanding.
Accurate perception of LiDAR point clouds is crucial for safe autonomous driving and advanced driver-assistance systems [9, 10, 11, 12, 13]. Traditional LiDAR point cloud models often depend on large annotated datasets, which are costly and time-consuming to create [14, 15]. To alleviate this challenge, research has explored semi-supervised [16, 17] and weakly-supervised [15, 18] methods. However, these approaches suffer from limited generalizability, especially when faced with diverse sensor configurations, e.g., different LiDAR beam numbers, camera placements, sampling rates, and potential sensor corruptions [19, 20, 10, 21, 22, 23, 24, 25, 26, 27]. This limitation poses a significant challenge for real-world scalability.
@@
A core innovation in our framework is the use of VFMs to generate semantically enriched superpixels from camera images, which are then aligned with LiDAR data to construct high-quality contrastive samples (see Figure 1). These semantic superpixels provide enhanced 2D-3D correspondences that capture object-level coherence, reducing the errors commonly associated with over-segmentation and “self-conflict” in contrastive learning [31]. This alignment significantly improves performance in downstream tasks, including 3D object detection and segmentation.
@@
As shown in Figure 2, compared to state-of-the-art methods such as SLidR [29] and ST-SLidR [30], our framework introduces significant advancements: i) the use of semantically rich superpixels to resolve the “self-conflict” issue in contrastive learning, ii) the creation of high-quality contrastive samples, which lead to faster and more stable convergence, and iii) reduced computational overhead due to the more efficient superpixel generation process.
@@ 2 Related Work
In this section, we provide a thorough literature review of relevant works in the areas of LiDAR-based autonomous driving scene understanding (Section 2.1), popular vision foundation models (Section 2.2), different 3D representation learning techniques (Section 2.3), and multi-dataset learning and utilization in the domain of 3D object detection and LiDAR semantic segmentation (Section 2.4).
@@ 2.1 LiDAR-Based Scene Understanding
For autonomous vehicles, accurate and dense 3D perception is crucial for safe navigation [11, 33]. Researchers have developed various methods for point cloud segmentation, including those based on raw points [34, 35, 36, 37, 38], range views [39, 40, 41, 42, 43, 44, 45], bird’s eye views [46, 47], voxels [48, 49, 50, 51], and multi-view fusion [52, 53, 54, 55, 56, 57]. Despite achieving significant advancements, these models typically rely on extensively annotated datasets, posing scalability issues [19]. To alleviate the annotation burden, recent studies have explored semi-supervised [16, 17, 58], weakly-supervised [15, 59, 60, 61, 18], and active learning [62, 63, 64] methods, as well as domain adaptation techniques [65, 66, 67, 68, 69, 70]. This work utilizes a self-supervised learning strategy, distilling knowledge from VFMs through camera-to-LiDAR associations, thereby eliminating the need for manual annotations during pretraining.
@@ 2.4 Multi-Dataset Utilization
Leveraging multiple datasets has emerged as a promising approach for improving the generalization of LiDAR-based models in autonomous driving [108]. Recent works like MDT3D [109] and MS3D++ [110] explored multi-source training for 3D object detection, while addressing challenges such as label space conflicts. Similarly, methods like COLA [111] and M3Net [112] utilized unified label spaces for semantic segmentation, demonstrating the advantages of multi-dataset learning. In this work, we extend these efforts by proposing a multi-source pretraining strategy that incorporates data from diverse LiDAR datasets, each with unique sensor setups and conditions [113, 114, 115, 13, 116]. By aligning 2D and 3D features across different sources, our approach improves cross-modal consistency and robustness. To the best of our knowledge, this is the first comprehensive study on multi-source pretraining for image-to-LiDAR representation learning, allowing our framework to generalize more effectively across varied driving environments and sensor configurations.
@@ 3 Image-to-LiDAR Data Pretraining
In this section, we elaborate on the technical details of the image-to-LiDAR pretraining. We first present the preliminaries (Section 3.1), followed by the introduction of superpixel-driven contrastive learning (Section 3.2), and conclude with the description of the contrastive learning objective (Section 3.3).
@@ 3.1 Problem Formulation
$[u_{i},v_{i},1]^{\text{T}}=\frac{1}{z_{i}}\times\Gamma_{K}\times\Gamma_{c\leftarrow l}\times[x_{i},y_{i},z_{i}]^{\text{T}}\penalty 10000\ ,$ (1)
@@ 4 LargeAD: A Scalable, Versatile, and Generalizable Framework
In this section, we present the technical details of the proposed framework. We begin by introducing our VFM-based superpixel generation approach (Section 4.1), followed by the explanation of semantic spatial consistency learning (Section 4.2). Next, we discuss the instance superpoint temporal consistency (Section 4.3), and conclude with the description of multi-source data pretraining (Section 4.4), culminating with a summary of the overall framework (Section 4.5).
@@ 4.1 Superpixel Generation from Foundation Models
Previous works have utilized SLIC [32] to group visually similar image regions into superpixels. However, SLIC often over-segments semantically coherent areas (see Figure 2), which poses challenges for contrastive learning, particularly due to the “self-conflict” phenomenon. This occurs when semantically similar superpixels are incorrectly treated as negative samples[120]. While [30] introduced a semantically tolerant loss to address this issue, the lack of high-level semantic understanding in SLIC exacerbates the difficulties in contrastive learning. To overcome these challenges, we generate semantic superpixels using vision foundation models (VFMs), which provide semantically rich superpixels and significantly improve the representation learning for both near and far points in the LiDAR point cloud (see Figure 5).
Instead of relying on low-level RGB features, our approach enhances superpixel generation by leveraging VFMs derived from large-scale pretrained image encoders [6, 8, 28, 7]. Unlike SLIC, VFMs capture high-level semantic information (as shown in Figure 2), allowing us to create more semantically meaningful superpixel sets, denoted as $\widehat{\Phi}_{\mathcal{S}}=\left\{\left\{\mathbf{s}_{m_{v}}\mid m_{v}=1,\ldots,M_{v}\right\}\mid M_{v}\ll M\right\}$. The generation process begins with the creation of semantic masks via prompts. By incorporating more abstract features, VFMs effectively address the “self-conflict” issue by grouping semantically similar regions more coherently, reducing the risk of misclassification during contrastive learning. As a result, the generated superpixels more accurately represent object semantics rather than just visual similarities. The corresponding superpoint set, $\widehat{\Phi}_{\mathcal{O}}=\left\{\left\{\mathbf{o}_{m_{v}}\mid m_{v}=1,\ldots,M_{v}\right\}\mid M_{v}\ll M\right\}$, is established using Eq. 1, ensuring proper alignment between the 2D image features and the 3D LiDAR point features.
@@ 4.2 Semantic Spatial Consistency Learning
To implement this framework, we use the same trainable LiDAR point cloud encoder $F_{\theta_{p}}$ and frozen image encoder $G_{\theta_{i}}$ as described earlier, extracting features from the LiDAR point cloud and 2D images, respectively. For the contrastive loss, we employ the projection heads $H_{\omega_{p}}$ and $H_{\omega_{i}}$ from Section 3.2, projecting both point and image features into a shared DD–dimensional embedding space. Unlike the low-level cues generated by SLIC, the superpixels produced by VFMs are enriched with semantic information, leading to more coherent and meaningful representations.
@@ 4.3 Instance Superpoint Temporal Consistency
Since the target embedding for all points within a segment in the current frame serves as the mean segment representation from the next frame, this loss will force points from a segment to converge to a mean representation while separating from other segments, implicitly clustering together points from the same instance. Figure 4 provides the positive feature correspondence in our contrastive learning framework. Furthermore, we swap $\widehat{\mathcal{F}}^{t}$ when generating the target mean embedding features to form a symmetric representation. In this way, the correspondence is encouraged from both $t\to{t+1}$ and ${t+1}\to t$, which leads to the following optimization objective: ${\mathcal{L}}^{\mathrm{tmp}}={\mathcal{L}}^{t\to{t+1}}+{\mathcal{L}}^{{t+1}\to t}$.
@@ 4.4 Multi-Source Data Pretraining
$\mathcal{L}^{\mathrm{cdp}}=-\frac{1}{M}{\sum_{i=1}^{M}\log\left[\frac{e^{\left(\left<\mathbf{k}^{(m)}_{i},\mathbf{k}^{(n)}_{i}\right>/\tau\right)}}{\sum_{j=1}^{M}e^{\left(\left<\mathbf{k}^{(m)}_{i},\mathbf{k}^{(n)}_{j}\right>/\tau\right)}}\right]}\penalty 10000\ .$ (7)
@@
The multi-source data pretraining leverages a variety of data sources to create a more resilient and flexible model. By addressing the substantial distributional differences between domains, feature normalization ensures consistency across diverse datasets, facilitating a more unified representation space. The cross-source contrastive loss further enhances adaptability, enabling it to transfer knowledge across multiple domains while maintaining domain-specific distinctions. As we will show in Section 5, this strategy results in improved generalization across different LiDAR sensors, leading to stronger performance on out-of-source downstream tasks.
Table 1: Benchmarking data pretraining methods pretrained on nuScenes [123] and fine-tuned on nuScenes [113], SemanticKITTI [114], and Waymo [115]. The evaluations are conducted on the official validation split of each dataset. Group (a): random initialization. Group (b): single-modality data pretraining. Group (c): image-to-LiDAR data pretraining by distilling ResNet [124]. Groups (d), (e), and (f): image-to-LiDAR data pretraining by distilling ViT [125] (Small, Medium, Large) from DINOv2 [73]. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores under each group are highlighted in bold and underline.
@@ 4.5 Overall Framework
$\mathcal{L}=\mathcal{L}^{\mathrm{vfm}}+\mathcal{L}^{\mathrm{tmp}}+\mathcal{L}^{\mathrm{p2s}}+\mathcal{L}^{\mathrm{cdp}}\penalty 10000\ .$ (8)
@@ 5 Experiments
In this section, we conduct extensive experiments to validate the effectiveness of the proposed LargeAD framework. We first introduce the datasets (Section 5.1) and describe the implementation details (Section 5.2). Next, we present a comparative study, demonstrating the superior performance of our framework compared to state-of-the-art methods across multiple tasks and datasets (Section 5.3). Finally, we provide an in-depth ablation study to analyze the impact of different components and data sources on the performance (Section 5.4).
@@ 5.1 Datasets
Table 2: Benchmarking data pretraining methods that are: (a) random initialized, (b) pretrained on SemanticKITTI [114], (c) pretrained on Waymo [115], and (d) pretrained based on our large-scale cross-sensor data pretraining pipeline. After pretraining, all methods are fine-tuned on SemanticKITTI [114], nuScenes [113], and Waymo [115], respectively, and evaluated on the official validation split of each dataset. LP denotes linear probing with frozen backbones. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
Table 3: Comparisons of different data pretraining methods that are: (a) random initialized, (b) pretrained on nuScenes [123], (c) pretrained on SemanticKITTI [114], (d) pretrained on Waymo Open [115], and (e) pretrained based on the proposed large-scale cross-sensor data pretraining pipeline. After pretraining, all methods are fine-tuned on different downstream datasets [15, 130, 131, 132, 133, 134, 69], respectively, and evaluated on the official validation split of each dataset. All mIoU scores are given in percentage (%). The best and second best scores are highlighted in bold and underline.
@@
Table 4: Robustness probing of data pretraining methods under eight out-of-distribution corruptions in the nuScenes-C dataset from the Robo3D benchmark [135]. The mCE score is the lower the better while mRR and mIoU scores are the higher the better. All mCE, mRR, and mIoU scores are given in percentage (%). Avg denotes the average mIoU scores of methods across all eight corruptions. The best and second best scores are highlighted in bold and underline.
@@
Figure 7: Qualitative results of object detection trained with $5\%$ labeled data. The first row shows the model trained with random initialization, while the second row displays results from our framework. The groundtruth / predicted results are highlighted with blue / red boxes, respectively. Best viewed in colors and zoomed-in for additional details.
5.2 Implementation Details
Evaluation Metrics. For linear probing and downstream tasks on LiDAR semantic segmentation, we report the Intersection-over-Union (IoU) for each semantic class and the mean IoU (mIoU) across all classes. For 3D object detection, we report the mean Average Precision (mAP) and nuScenes Detection Score (NDS). For panoptic LiDAR segmentation, we follow the Panoptic-nuScenes protocol [123] and report the Panoptic Quality (PQ), Recognition Quality (RQ), and Segmentation Quality (SQ). For robustness evaluation, we adopt the Robo3D benchmark [135] and report the mean Corruption Error (mCE) and mean Resilience Rate (mRR) scores. The baseline model used for these evaluations is the MinkUNet implementation in [53].
5.3 Comparative Study
Comparisons to State-of-the-Arts. We compare the proposed LargeAD against random initialization and a total of eleven state-of-the-art pretraining techniques using both linear probing (LP) and few-shot fine-tuning protocols on nuScenes [123], as presented in Table 1. The results demonstrate the significant impact of pretraining on downstream task performance, particularly in low-data regimes such as $1\%$, $5\%$, and $10\%$ fine-tuning budgets. When distilling the knowledge from ResNet, ViT-S, ViT-B, and ViT-L, our framework achieves mIoU scores of $46.13\%$, $46.58\%$, $47.84\%$, and $48.71\%$ under the LP setting, outperforming the previous best models [30, 31, 127, 129] by large margins. Moreover, our framework consistently delivers the highest performance across almost all fine-tuning tasks on nuScenes, highlighting the effectiveness of VFM-assisted contrastive learning, the incorporation of spatial-temporal consistency regularization, and the combination of multi-source data pretraining.
Downstream Generalization. To thoroughly evaluate the generalization capability of our method, we perform experiments across a total of nine autonomous driving datasets, with the results summarized in Table 1 (for SemanticKITTI and Waymo Open) and Table 3 (for the other seven datasets). Each dataset presents distinct challenges, including variations in sensor types, acquisition environments, scale, and data fidelity, making this a rigorous assessment of model generalization. Our framework achieves $51.68\%$ and $52.68\%$ mIoU scores on SemanticKITTI and Waymo Open when distilling from ViT-L, setting new states for these benchmarks. We also surpass SLidR [29] and Seal [31] on the other seven datasets, as shown in Table 3. The results consistently show that the proposed method outperforms existing state-of-the-art approaches on all evaluated datasets. These outcomes underline the robustness and adaptability of our proposed approach to a wide range of real-world automotive perception tasks.
Table 5: Comparisons of state-of-the-art pretraining methods pretrained and fine-tuned on nuScenes [113] using the specified data proportions. All methods employ CenterPoint [137] or SECOND [138] as the 3D object detection backbones. All sores are given in percentage (%). The best and second best scores are highlighted in bold and underline.